# A Wireless AI-Generated Content (AIGC) Provisioning Framework Empowered by Semantic Communication

### v2

@@ III-B Decoder Design at Local Receiver
1) Channel decoder module: The received binary bits are recovered as semantic information by a traditional channel decoder. The decoded information can be represented as
${z}^{\prime}=\mathcal{C}^{-1}({y}^{\prime}),$ (6)
where $\mathcal{C}^{-1}$ is the channel decoder. Therefore, $\epsilon^{\cal C}=z^{\prime}-z,\epsilon^{\cal C}\sim{\cal N}(0,\sigma^{2}I)$ is the semantic noise with a variance matrix $\sigma^{2}I$ caused by the physical channel noise $n$.
2) Fine-tuning module: This is a lightweight version of the semantic generation module, which is used for fine-tuning decoded image semantic information [27]. The reverse diffusion process of fine-tuning module is as follows
$z_{\bar{\theta}}=\mathcal{F}_{2}({z}^{\prime},\bar{T},z_{\scriptscriptstyle t\!e\!x\!t};{\bar{\theta}}),$ (7)
where $\mathcal{F}_{2}(\cdot;\bar{\theta})$ is the UNet with learnable parameter $\bar{\theta}$, $\bar{T}$ is the number of reverse diffusion steps for the fine-tuning module. Nevertheless, unlike the semantic generation module, the fine-tuning module considers semantic noise $\epsilon^{\cal C}$ when training the UNet. We theoretically prove that semantic noise $\epsilon^{\cal C}$ can be denoised according to the modified reverse diffusion process, as shown in the Proposition. 1.
@@
Given the semantic noise $\epsilon^{\cal C}\sim{\cal N}(0,\sigma^{2}I)$, the reverse diffusion process from denoising step $t$ to $t-1$ in the fine-tuning module is:
$z_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(z_{t}-C(\alpha,\sigma,t)\epsilon_{\theta}(z_{t},t,z_{\scriptscriptstyle t\!e\!x\!t})\right)+\bar{\sigma}_{t}\epsilon,$ (8)
$C(\alpha,\sigma,t)=\frac{\left(1-\alpha_{t}+\sigma_{t,t-1}^{2}\right)\left(\sqrt{1-\dot{\alpha}_{t}}-\sigma_{t}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}},$
where $\epsilon_{\theta}(z_{t},t,z_{\scriptscriptstyle t\!e\!x\!t})$ is the noise predicted from semantic information $z_{t}$, and $C(\alpha,\sigma,t)$ is a set of time related constants. Additionally, $\sigma_{t}$, $\sigma_{t-1}$, $\sigma_{t,t-1}$, and $\bar{\sigma}_{t}$ are time-dependent standard deviations for the random sample of Gaussian noise.
@@
$z_{t}$ $=\sqrt{\alpha_{t}}z_{t-1}+\sqrt{1-\alpha_{t}}\epsilon+\epsilon^{\cal C}$ (9)
$=\sqrt{\dot{\alpha_{t}}}z_{0}+\sqrt{1-\dot{\alpha_{t}}}\epsilon+\epsilon^{\cal C},$
$\epsilon^{\cal C}\sim{\cal N}(0,\sigma^{2}I),\epsilon\sim{\cal N}(0,I).$
However, for model training, we consider that the semantic noise $\epsilon^{\cal C}$ is gradually added in the diffusion steps $\{1,\cdots,t,\cdots,\bar{T}\}$, then the diffusion process can be formulated as
$z_{t}$ $=\sqrt{\alpha_{t}}z_{t-1}+\sqrt{1-\alpha_{t}}\epsilon+\sqrt{\gamma_{t}}\epsilon^{\cal C}$ (10)
$=\sqrt{\dot{\alpha}_{t}}z_{0}+\sqrt{1-\dot{\alpha}_{t}}\epsilon+\sigma\sum_{t=1}^{\bar{T}}\prod_{j=t+1}^{\bar{T}-1}\sqrt{\gamma_{t}}\sqrt{\alpha_{j}}\epsilon,$
@@
The aim of fine-tuning module is to generate image semantic information $z_{0}$ from received semantic information $z_{t}$ via the reverse diffusion process. Therefore, it is important to predict the distribution probability of the semantic information in the previous step $z_{t-1}$ with the present information $z_{t}$, i.e., $q(z_{t-1}|z_{t},z_{0})$. According to the Bayes’ theorem, we have
$q(z_{t-1}|z_{t},z_{0})=q(z_{t}|z_{t-1},z_{0})\frac{q(z_{t-1}|z_{0})}{q(z_{t}|z_{0})},$ (11)
$q(z_{t}|z_{0})\sim{\cal N}(\sqrt{\dot{\alpha}_{t}}z_{0},1-\dot{\alpha}_{t}+{\sigma}_{t}^{2}),$
$q(z_{t-1}|z_{0})\sim{\cal N}(\sqrt{\dot{\alpha}_{t-1}}z_{0},1-\dot{\alpha}_{t-1}+{\sigma}_{t-1}^{2}),$
$q(z_{t}|z_{t-1},z_{0})\sim{\cal N}(\sqrt{\dot{\alpha}_{t}}z_{t-1},1-\alpha_{t}+{\sigma}_{t,t-1}^{2}).$
@@
Here, noise variances ${\sigma}_{t}^{2}$, ${\sigma}_{t-1}^{2}$, and ${\sigma}_{t,t-1}^{2}$ can be determined according to $\sigma\prod_{j=t+1}^{\bar{T}-1}\sqrt{\gamma_{t}}\sqrt{\alpha_{j}}$. Then, with the Gaussian distribution formula $f(x)=\frac{1}{\sigma\sqrt{2\pi}}exp\left(-\frac{1}{2}\left(\frac{x-\mu}{\sigma}\right)^{2}\right)$, we have
$q(z_{t-1|z_{t},z_{0}})\varpropto$ (12)
$\frac{1}{\sigma\sqrt{2\pi}}exp\left(-\frac{1}{2}\left(\frac{\left(z_{t}-\sqrt{\dot{\alpha}_{t}}z_{t-1}\right)^{2}}{1-\alpha_{t}+{\sigma}_{t,t-1}^{2}}\right.\right.$
$\left.\left.+\frac{\left(z_{t-1}-\sqrt{\dot{\alpha}_{t-1}}z_{0}\right)^{2}}{1-\dot{\alpha}_{t-1}+{\sigma}_{t-1}^{2}}-\frac{\left(z_{t}-\sqrt{\dot{\alpha}_{t}}z_{0}\right)^{2}}{1-\dot{\alpha}_{t}+{\sigma}_{t}^{2}}\right)\right).$
@@
Since the purpose of the reverse diffusion process is to predict the semantic information in the previous step, while $z_{t}$ and $z_{0}$ are known semantic information, we have
$q(z_{t-1|z_{t},z_{0}})\varpropto\frac{1}{\sigma\sqrt{2\pi}}\times$ (13)
$exp\left(-\frac{1}{2}\left(\left(\frac{\alpha_{t}}{1-\alpha_{t}+{\sigma}_{t,t-1}^{2}}+\frac{1}{1-\dot{\alpha}_{t-1}+{\sigma}_{t-1}^{2}}\right)z_{t-1}^{2}\right.\right.$
$\left.\left.-2\!\left(\frac{\sqrt{\alpha_{t}}z_{t}}{1\!-\!\alpha_{t}\!+\!{\sigma}_{t,t-1}^{2}}\!+\!\frac{\sqrt{\dot{\alpha}_{t-1}}z_{0}}{1\!-\!\dot{\alpha}_{t-1}\!+\!{\sigma}_{t-1}^{2}}\right)\!z_{t-1}\!+\!C\left(z_{t},z_{0}\right)\!\right)\right).$
@@
As Gaussian distribution formula can be transformed as $f(x)=\frac{1}{\sigma\sqrt{2\pi}}exp\left(-\frac{1}{2}\left(\frac{1}{\sigma^{2}}x^{2}-\frac{2\mu}{\sigma^{2}}x+\frac{\mu^{2}}{\sigma^{2}}\right)\right)$, we can calculate the mean of the distribution as
$\mu_{\theta}(z_{t},z_{0})=\frac{\sqrt{\alpha_{t}}\left(1-\dot{\alpha}_{t-1}+\sigma_{t-1}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}}z_{t}+$ (14)
$\frac{\sqrt{\dot{\alpha}_{t-1}}\left(1-\alpha_{t}+\sigma_{t,t-1}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}}z_{0}.$
@@
Then, (10) can be reversed as $z_{0}=\frac{1}{\sqrt{\dot{\alpha}_{t}}}\left(z_{t}-(\sqrt{1-\dot{\alpha}_{t}}-\sigma_{t}^{2})\epsilon_{t}\right),\epsilon_{t}\sim{\cal N}(0,I)$, thus we have
@@
$\mu_{\theta}(z_{t},t)\!=\!\frac{1}{\sqrt{\alpha_{t}}}\left(z_{t}\!-\!\frac{\left(1\!-\!\alpha_{t}\!+\!\sigma_{t,t-1}^{2}\right)\left(\sqrt{1\!-\!\dot{\alpha}_{t}}\!-\!\sigma_{t}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}}\epsilon_{t}\right).$ (15)
Therefore, to sample $z_{t-1}\sim p_{\bar{\theta}}(z_{t-1}|z_{t})$ is to compute $z_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(z_{t}-C(\alpha,\sigma,t)\epsilon_{\theta}(x_{t},t)\right)+\bar{\sigma}_{t}\epsilon$, we prove Proposition 1. ∎
@@
$\bar{s}=\mathcal{D}(z_{\bar{\theta}}|z_{\scriptscriptstyle t\!e\!x\!t};\bar{\varphi}),$ (16)
where $\mathcal{D}(\cdot;\bar{\varphi})$ is the autoencoder decoder with learnable parameter $\bar{\varphi}$.
III-C Training Algorithm
The training of image process modules is the major part of generating a high-quality image from text. Specifically, we first introduce the joint training of the image semantic extraction module and autoencoder decoder module. Then, we present the training of diffusion models in the semantic generation module and the fine-tuning module. The text-extracting module can use a pre-trained CLIP model since it is not a main focus.
The image semantic extraction module and the autoencoder decoder module are jointly trained according to the loss function of the VAE encoder-decoder [28]. In the consideration of semantic similarity and reconstruction likelihood, the loss function is
$loss_{1}(\varphi,\bar{\varphi})=$ (17)
$-\frac{1}{2m}\sum_{i=1}^{M}\sum_{j=1}^{K}\left(1+\log\sigma_{j}^{2}-\sigma_{j}^{2}-\mu_{j}^{2}\right)+\frac{1}{m}\sum_{i=1}^{M}(s_{i}-\bar{s}_{i})^{2}$
$=-D_{\scriptscriptstyle K\!L}\left(\tilde{q}_{\varphi}\left(z|s\right)||\tilde{p}\left(z\right)\right)+\mathbb{E}_{z\sim\tilde{q}_{\varphi}\left(s|z\right)}\left[log\tilde{p}_{\bar{\varphi}}\left(s|z\right)\right],$
where $i\in\{1,\cdots,M\}$ is the $i^{th}$ sample in a dataset with a total size $M$, $K$ is the set size of means $\mu_{j}$ and standard deviations $\sigma_{j}$ generated by the VAE encoder, $\tilde{q}_{\varphi}(z|s)$ is the distribution of the VAE encoder output semantic information, $\tilde{p}(z)$ the distribution of a real image’s semantic information, $\tilde{p}_{\bar{\varphi}}(s|z)$ is the distribution of the VAE decoder output, and $D_{\scriptscriptstyle K\!L}$ is the cross-entropy.

### v3

@@ 3.2 Decoder Design at Local Receiver
1) Channel decoder module: The received binary bits are recovered as semantic information by a traditional channel decoder.
The decoded information can be represented as
$\mathbf{z}^{\prime}=\mathcal{C}^{-1}(\mathbf{Y}^{\prime}),$ (6)
where $\mathcal{C}^{-1}$ is the channel decoder. Therefore, $\bm{\epsilon}^{\cal C}=\mathbf{z}^{\prime}-\mathbf{z},\bm{\epsilon}^{\cal C}\sim{\cal N}(0,\sigma^{2}\bm{I})$ is the semantic noise with a variance matrix $\sigma^{2}\bm{I}$ caused by the physical channel noise $\mathbf{N}$.
2) Fine-tuning module: This is a lightweight version of the semantic generation module, which is used for fine-tuning decoded image semantic information. The reverse diffusion process of fine-tuning module is as follows
$\mathbf{z}_{\bar{\bm{\theta}}}=\mathcal{F}_{2}(\mathbf{z}^{\prime},\bar{T},\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t};{\bar{\bm{\theta}}}),$ (7)
where $\mathcal{F}_{2}(\cdot;\bar{\bm{\theta}})$ is the UNet with learnable parameter $\bar{\bm{\theta}}$, $\bar{T}$ is the number of reverse diffusion steps for the fine-tuning module. Nevertheless, unlike the semantic generation module, the fine-tuning module considers semantic noise $\bm{\epsilon}^{\cal C}$ when training the UNet. We theoretically prove that semantic noise $\bm{\epsilon}^{\cal C}$ can be denoised according to the modified reverse diffusion process, as shown in the Proposition. 1.
@@
Assume semantic noise $\bm{\epsilon}^{\cal C}\sim{\cal N}(0,\sigma^{2}\bm{I})$ is gradually added to semantic information, the reverse diffusion process from denoising step $t$ to $t-1$ in the fine-tuning module is
$\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-C(\alpha,\sigma,t)\bm{\epsilon}_{\bm{\theta}}(\mathbf{z}_{t},t,\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t})\right)+\bar{\sigma}_{t}\bm{\epsilon},$ (8)
$where$ $C(\alpha,\sigma,t)=\frac{\left(1-\alpha_{t}+\sigma_{t,t-1}^{2}\right)\left(\sqrt{1-\dot{\alpha}_{t}}-\sigma_{t}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}}$
is a set of time related constants, and $\bm{\epsilon}_{\bm{\theta}}(\mathbf{z}_{t},t,\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t})$ is the noise predicted from semantic information $\mathbf{z}_{t}$. Additionally, $\sigma_{t}$, $\sigma_{t-1}$, $\sigma_{t,t-1}$, and $\bar{\sigma}_{t}$ are time-dependent standard deviations for the random sample of Gaussian noise.
@@
$\mathbf{z}_{t}$ $=\sqrt{\alpha_{t}}\mathbf{z}_{t-1}+\sqrt{1-\alpha_{t}}\bm{\epsilon}+\bm{\epsilon}^{\cal C}$ (9)
$=\sqrt{\dot{\alpha_{t}}}\mathbf{z}_{0}+\sqrt{1-\dot{\alpha_{t}}}\bm{\epsilon}+\bm{\epsilon}^{\cal C},$
$\bm{\epsilon}^{\cal C}\sim{\cal N}(0,\sigma^{2}\bm{I}),\bm{\epsilon}\sim{\cal N}(0,\bm{I}).$
Here, semantic noise $\bm{\epsilon}^{\cal C}$ is added after transmission at one time, while in general diffusion, Gaussian noise is gradually added and denoised through multiple steps. If denoising all semantic noise within a single step, minor noise prediction errors can lead to semantic information being lost or distorted. Therefore, it is necessary to simulate the general diffusion process to gradually filter channel-caused semantic noise $\bm{\epsilon}^{\cal C}$. We assume semantic noise is gradually added during transmission, thus vice it can be gradually removed. With diffusion steps $\{1,\cdots,t,\cdots,\bar{T}\}$, the diffusion process can be formulated as
$\mathbf{z}_{t}$ $=\sqrt{\alpha_{t}}\mathbf{z}_{t-1}+\sqrt{1-\alpha_{t}}\bm{\epsilon}+\sqrt{\gamma_{t}}\bm{\epsilon}^{\cal C}$ (10)
$=\sqrt{\dot{\alpha}_{t}}\mathbf{z}_{0}+\sqrt{1-\dot{\alpha}_{t}}\bm{\epsilon}+\sigma\sum_{t=1}^{\bar{T}}\prod_{j=t+1}^{\bar{T}-1}\sqrt{\gamma_{t}}\sqrt{\alpha_{j}}\bm{\epsilon},$
@@
The aim of fine-tuning module is to generate image semantic information $\mathbf{z}_{0}$ from received semantic information $\mathbf{z}_{t}$ via the reverse diffusion process. Therefore, it is important to predict the distribution probability of the semantic information in the previous step $\mathbf{z}_{t-1}$ with the present information $\mathbf{z}_{t}$, i.e., $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0})$. According to the Bayes’ theorem, we have
$q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0})=q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{0})\frac{q(\mathbf{z}_{t-1}|\mathbf{z}_{0})}{q(\mathbf{z}_{t}|\mathbf{z}_{0})}.\$ (11)
The distributions are Gaussian, and according to equation (10), we have
$q(\mathbf{z}_{t}|\mathbf{z}_{0})\sim{\cal N}(\sqrt{\dot{\alpha}_{t}}\mathbf{z}_{0},1-\dot{\alpha}_{t}+{\sigma}_{t}^{2}),$ (12)
$q(\mathbf{z}_{t-1}|\mathbf{z}_{0})\sim{\cal N}(\sqrt{\dot{\alpha}_{t-1}}\mathbf{z}_{0},1-\dot{\alpha}_{t-1}+{\sigma}_{t-1}^{2}),$
$q(\mathbf{z}_{t}|\mathbf{z}_{t-1},\mathbf{z}_{0})\sim{\cal N}(\sqrt{\dot{\alpha}_{t}}\mathbf{z}_{t-1},1-\alpha_{t}+{\sigma}_{t,t-1}^{2}).$
@@
Here, noise variances ${\sigma}_{t}^{2}$, ${\sigma}_{t-1}^{2}$, and ${\sigma}_{t,t-1}^{2}$ can be determined according to $\sigma\prod_{j=t+1}^{\bar{T}-1}\sqrt{\gamma_{t}}\sqrt{\alpha_{j}}$. Then, with the Gaussian distribution formula $f(x)=\frac{1}{\sigma\sqrt{2\pi}}exp\left(-\frac{1}{2}\left(\frac{x-\mu}{\sigma}\right)^{2}\right)$, and as per equation (11), we have
$q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0})\varpropto$ (13)
$\frac{1}{\sigma\sqrt{2\pi}}exp\left(-\frac{1}{2}\left(\frac{\left(\mathbf{z}_{t}-\sqrt{\dot{\alpha}_{t}}\mathbf{z}_{t-1}\right)^{2}}{1-\alpha_{t}+{\sigma}_{t,t-1}^{2}}\right.\right.$
$\left.\left.+\frac{\left(\mathbf{z}_{t-1}-\sqrt{\dot{\alpha}_{t-1}}\mathbf{z}_{0}\right)^{2}}{1-\dot{\alpha}_{t-1}+{\sigma}_{t-1}^{2}}-\frac{\left(\mathbf{z}_{t}-\sqrt{\dot{\alpha}_{t}}\mathbf{z}_{0}\right)^{2}}{1-\dot{\alpha}_{t}+{\sigma}_{t}^{2}}\right)\right).$
@@
Since the purpose of the reverse diffusion process is to predict the semantic information in the previous step, while $\mathbf{z}_{t}$ and $\mathbf{z}_{0}$ are known semantic information, we have
$q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0})\varpropto\frac{1}{\sigma\sqrt{2\pi}}\times$ (14)
$exp\left(-\frac{1}{2}\left(\left(\frac{\alpha_{t}}{1-\alpha_{t}+{\sigma}_{t,t-1}^{2}}+\frac{1}{1-\dot{\alpha}_{t-1}+{\sigma}_{t-1}^{2}}\right)\mathbf{z}_{t-1}^{2}\right.\right.$
$\left.\left.-2\!\left(\frac{\sqrt{\alpha_{t}}\mathbf{z}_{t}}{1\!-\!\alpha_{t}\!+\!{\sigma}_{t,t-1}^{2}}\!+\!\frac{\sqrt{\dot{\alpha}_{t-1}}\mathbf{z}_{0}}{1\!-\!\dot{\alpha}_{t-1}\!+\!{\sigma}_{t-1}^{2}}\right)\!\mathbf{z}_{t-1}\!+\!C\left(\mathbf{z}_{t},\mathbf{z}_{0}\right)\!\right)\right).$
@@
As Gaussian distribution formula can be transformed as $f(x)=\frac{1}{\sigma\sqrt{2\pi}}exp\left(-\frac{1}{2}\left(\frac{1}{\sigma^{2}}x^{2}-\frac{2\mu}{\sigma^{2}}x+\frac{\mu^{2}}{\sigma^{2}}\right)\right)$, we have $\frac{\mu}{\sigma^{2}}=\left(\frac{\sqrt{\alpha_{t}}\mathbf{z}_{t}}{1\!-\!\alpha_{t}\!+\!{\sigma}_{t,t-1}^{2}}\!+\!\frac{\sqrt{\dot{\alpha}_{t-1}}\mathbf{z}_{0}}{1\!-\!\dot{\alpha}_{t-1}\!+\!{\sigma}_{t-1}^{2}}\right)$ and $\frac{1}{\sigma^{2}}=\left(\frac{\alpha_{t}}{1-\alpha_{t}+{\sigma}_{t,t-1}^{2}}+\frac{1}{1-\dot{\alpha}_{t-1}+{\sigma}_{t-1}^{2}}\right)$. Then, we can calculate the mean of the distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0})$ as
$\bar{\mu}(\mathbf{z}_{t},\mathbf{z}_{0})\!=\frac{\sqrt{\alpha_{t}}\left(1-\dot{\alpha}_{t-1}+\sigma_{t-1}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}}\mathbf{z}_{t}+$ (15)
$\frac{\sqrt{\dot{\alpha}_{t-1}}\left(1-\alpha_{t}+\sigma_{t,t-1}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}}\mathbf{z}_{0}.$
@@
Then, since (10) can be reversed as $\mathbf{z}_{0}=\frac{1}{\sqrt{\dot{\alpha}_{t}}}\left(\mathbf{z}_{t}-(\sqrt{1-\dot{\alpha}_{t}}-\sigma_{t}^{2})\bm{\epsilon}_{t}\right),\bm{\epsilon}_{t}\sim{\cal N}(0,\bm{I})$, thus we have
$\bar{\mu}(\mathbf{z}_{t},t)$ $=\frac{\sqrt{\alpha_{t}}\left(1-\dot{\alpha}_{t-1}+\sigma_{t-1}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}}\mathbf{z}_{t}+$ (16)
$\frac{1-\alpha_{t}+\sigma_{t,t-1}^{2}}{{\sqrt{\alpha}_{t}}\left(1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}\right)}\mathbf{z}_{t}-$
$\frac{\left(1-\alpha_{t}+\sigma_{t,t-1}^{2}\right)\left(\sqrt{1-\dot{\alpha}_{t}}-\sigma_{t}^{2}\right)}{{\sqrt{\alpha}_{t}}\left(1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}\right)}\bm{\epsilon}_{t}$
@@
$=\!\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}\!-\!\frac{\left(1\!-\!\alpha_{t}\!+\!\sigma_{t,t-1}^{2}\right)\left(\sqrt{1\!-\!\dot{\alpha}_{t}}\!-\!\sigma_{t}^{2}\right)}{1-\dot{\alpha}_{t}+\sigma_{t-1}^{2}\alpha_{t}+\sigma_{t,t-1}^{2}}\bm{\epsilon}_{t}\right).$
Therefore, we have $\mu_{\bm{\theta}}(\mathbf{z}_{t},t)$ as the mean of distribution $q(\mathbf{z}_{t-1}|\mathbf{z}_{t},\mathbf{z}_{0})$. Using a neural network to estimate $\bm{\epsilon}_{t}$, and according to the reparameterization trick, we have $\mathbf{z}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{z}_{t}-C(\alpha,\sigma,t)\bm{\epsilon}_{\bm{\theta}}(x_{t},t)\right)+\bar{\sigma}_{t}\bm{\epsilon}$.
@@
$\bar{s}=\mathcal{D}(\mathbf{z}_{\bar{\bm{\theta}}}|\mathbf{z}_{\scriptscriptstyle t\!e\!x\!t};\bar{\bm{\varphi}}),$ (17)
where $\mathcal{D}(\cdot;\bar{\bm{\varphi}})$ is the autoencoder decoder with learnable parameter $\bar{\bm{\varphi}}$.
3.3 Training Algorithm
The training of image process modules is the key to generating a high-quality image from text. Specifically, the image semantic extraction module and the autoencoder decoder module are jointly trained according to the loss function of the VAE encoder-decoder [26, 31] with a dataset $\mathbf{S}=\{\mathbf{s}^{(1)},\cdots,\mathbf{s}^{(i)},\cdots\mathbf{s}^{(M)}\}$, $M$ is the dataset size. In the consideration of semantic similarity and reconstruction likelihood of each image $\mathbf{s}^{(i)}$, the loss function is
$loss_{1}(\bm{\varphi},\bar{\bm{\varphi}})=$ $\frac{1}{2m}\!\sum_{i=1}^{M}\sum_{j=1}^{K}\!\left(1\!+\!\log\sigma_{i,j}^{2}\!-\!\sigma_{i,j}^{2}\!-\!\mu_{i,j}^{2}\right)$ (18)
$\!+\!\frac{1}{m}\sum_{i=1}^{M}\left(\|\mathbf{s}^{(i)}\!-\!\bar{\mathbf{s}}^{(i)}\|_{2}^{2}\right)\!,$
where $K$ is the set size of means $\mu_{i,j}\in\bm{\mu}_{i}$, and standard deviations $\sigma_{i,j}\in\bm{\sigma}_{i}$ generated for datapoint $\mathbf{s}^{(i)}$ by the VAE encoder. Here, the semantic similarity term $\sum_{j=1}^{K}\!\left(1\!+\!\log\sigma_{i,j}^{2}\!-\!\sigma_{i,j}^{2}\!-\!\mu_{i,j}^{2}\right)$ is the KL divergence between the approximate Gaussian postier $\mathcal{N}(\bm{\mu}_{i},\bm{\sigma}_{i}^{2})$ and standard normal prior $\mathcal{N}(0,\bm{I})$. Meanwhile, $\left(\|\mathbf{s}_{i}\!-\!\bar{\mathbf{s}}_{i}\|_{2}^{2}\right)$ is the restructuration likelihood term, which measures the difference between the original image $\bar{\mathbf{s}}_{i}$ and the reconstructed image $\bar{\mathbf{s}}_{i}$.